Circular No. 8982

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2008 K3–K11 AND C/2008 L4 (SOHO) Further to IAUC 8981, additional near-sun presumed comets have been

found on SOHO website images — all Kreutz sungrazers except for C/2008 K7 (Meyer group), which was tiny, stellar in appearance, and faint (mag ∼ 7.5–8), and C/2008 K10, which was small and stellar in appearance (mag 7–7.5, brightening as it left the C2 field-of-view but never appeared in C3). C/2008 K10 has been identified by R. Kracht with C/1999 X3 (cf. IAUC 8735; not reported until 2006 and being near the limit of visibility in C2 images) and C/2004 E2 (cf. IAUC 8365), an identification confirmed by B. G. Marsden (with linked orbits and residuals published on MPC 2008-S49; for epoch 2008 May 14.0 TT, T = 2008 May 31.334 TT, q = 0.04797 AU, e = 0.98162, ω = 353^{o}.611, Ω = 323^{o}.655, i = 6^{o}.259, equinox 2000.0); Marsden notes that close approaches occurred to the earth on 2000 Jan. 13 (Δ = 0.058 AU) and Mars on 2004 May 19 (Δ_{M} = 0.032 AU). K. Battams writes that C/2008 K3 was tiny, extremely faint (mag ∼ 8.5), and stellar in appearance. C/2008 K4 was one of the brightest comets seen by SOHO (saturating slightly in both LASCO cameras, with estimated peak magnitude ∼ 1–2), appearing as a bright teardrop in C3 images and having a partly-forked, “thick” tail ∼ 0^{o}.5 long in C2 images; both SECCHI COR cameras also imaged C/2008 K4, with the tail appearing somewhat thinner in the COR2A images and showing the slight “fork” in the COR2B images. C/2008 K5 and C/2008 K7 were tiny and stellar in appearance (mag ∼ 7.5–8). C/2008 K6 was tiny and stellar in appearance (mag ∼ 7) in C3 images, and elongated and very diffuse in C2 images. C/2008 K8 was very diffuse (mag ∼ 8). C/2008 K9 and C/2008 K11 were both small and very diffuse (mag ∼ 8–8.5, the former being perhaps slightly fainter than the latter). C/2008 L4 was small and slightly diffuse (mag ∼ 7–7.5).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 K3 May 17.858 3^{h}41.^{m}0 +17^{o}51′ C2  MK 2008-M13
C/2008 K4  21.988 4 08.3 +16 12 C3/2 RM 2008-M13
C/2008 K5 23.579 4 02.4 +18 52 C2 MK 2008-M13
C/2008 K6  25.513 4 10.2 +18 08 C3/2 HS 2008-O15
C/2008 K7  25.913 4 13.3 +22 23 C2  JR 2008-O15
C/2008 K8  28.079 4 20.9 +19 43 C2  JR 2008-O15
C/2008 K9  28.704 4 23.1 +19 47 C2  GP 2008-O15
C/2008 K10  30.881 4 27.8 +21 22 C2  RK 2008-O16
C/2008 K11  31.038 4 31.8 +20 05 C2  GS 2008-O16
C/2008 L4  June 2.371 4 39.4 +20 18 C2  MK 2008-O16
2008 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green